GOOD DAY COULD YOU PLEASE HELP ME: A father is twice the age of his son. In 10 years time the father would have been 52 years old, how old would his son be by then?
Answer by richwmiller(17219)

52-10=42 =f
s=f/2=42/2=21
s+10=31
the father is 42 now and the son is 21 now
when the father is 52 the son will be 31